Garcia, Inc. uses a job-order costing system for its products, which pass from the Machining Department, to the Assembly Depa4. Determine whether overhead was under- or overapplied during the year in the Assembly Department. Coverapplied overhead6. How much overhead would have been charged to the companys Work-in-process account during the year? Work-in-process
Answer #1
2
Machining Assembly
a Budgeted Manufacturing Overhead 4,000,000 3,080,000
b Budgeted Machine Hours 400,000
c Budgeted Direct labor cost           5,600,000
d Overhead Rate of
Machining =(a/b)
Assembly= (a/c)
10                      0.55
Machining Assembly Total
a Direct Material 23,500 6,800 30,300
b Direct Labor 27,800 58600 86,400
c Overhead cost for Machining 3700 3,700
d Overhead cost for Assembly           32,230.00 32,230
e Total cost of work in Process 55,000           97,630.00 152,630
Total cost of work in Process 152,630
Calculation of Overhead cost
overhead cost of Machining = Machine hour * Overhead rate
Machining= 370*10= 3700
Overhead cost of Assembly = direct labor cost * overhead rate
Overhead cost of Assembly = 58600*0.55=32230
4
Machining Assembly
a applied overhead cost
Machining = Overhead rate * actual Machine hours
= 10*425,000
Assembly = overhead rate * actual direct labor cost = 0.55*5780000
          4,250,000.00    3,179,000.00
b actual overhead cost           4,270,000.00     3,030,000.00
c Over or under applied overhead (a-b)                  (20,000.00)         149,000.00
Matching department under - here applied because applied Manufacturing overhead cost is lower than actual overhead cost
Assembly Department over- here applied, applied Manufacturing overhead cost is higher than actual over head cost
6 Overhead Charged to Work in process during the year
Machining Assembly Total
applied Manufacturing Overhead           4,250,000.00    3,179,000.00    7,429,000.00
Overhead Charged to Work in process during the year =7,429,000 (sum of Both department)